OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (8, 12) and has a slope of −2? y = ? x + ? help

OpenStudy (anonymous):

y=+-2=28?

OpenStudy (zpupster):

you have x1 and y1 (x1,y1)or here (8,12) we use the point slope-form to get our slope-intercept form point slope-form y-y1=m(x-x1) so we plug in the values y-12= -2 (x- 8) multiply thru

OpenStudy (anonymous):

kay i have it :) thanks

OpenStudy (zpupster):

good job!